OpenStudy (anonymous):

what is the rate of consumption of each reactant when the rate of production of ammonia is 4.0 x10 ^ -3 mol/L/s N2 (g) + 3H2 (g) ---> 2NH3 (g)

OpenStudy (aaronq):

do you know how to write the differential rates?

OpenStudy (aaronq):

For example you have: \(3A⇌2B\) the production of B is: \(rate=\dfrac{1}{2}\dfrac{d[B]}{dt}\) which is equal to the consumption of A: \(rate=-\dfrac{1}{3}\dfrac{d[A]}{dt}\) they're equal to each other so \(rate=-\dfrac{1}{3}\dfrac{d[A]}{dt}=\dfrac{1}{2}\dfrac{d[B]}{dt}\) Note that since A is being consumed, it is a negative rate.